# HoneySat: A Network-based Satellite Honeypot Framework

### v1

@@ 2.5 Small Satellite Protocol Ecosystems
SmallSat missions can often be categorized into the following categories by the adoption of protocols and their corresponding philosophies.
Cubesat Space Protocol. The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for Small missions [63]. There are not a lot of choices left to the operator/developer and there are two vendors mainly offering components using this protocol and some that offer compatibility solutions [23]. If a mission built on CSP is expanded by a commercial subsystem it will plug in without issue after configuration.
@@
CSP is implemented as an open-source C library called libCSP [63]. CSP follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), CAN (Controller Area Network), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [72].
In CSP the ground segment interfaces and satellite subsystems are part of a CSP network. Sending TC is as simple as sending a CSP packet with an address corresponding to a node that is part of the satellite. The configured static routing tables of the nodes on the network will then cause the packet to be passed to the ground station and transmitted to the satellite over RF where a system on board will route the packet to its destination. In case the satellite is not currently doing a pass, the packet is dropped.
The straightforwardness of CSP makes it particularly interesting for the honeypot use case as the internal structures of any CSP-based mission will look very similar, so distinguishing two missions is difficult as they will use the protocols in a very similar way.
@@
There are also default services running on standardized ports that are enabled by default when building libCSP [63], for example for network diagnostic (ping) or basic operations (rebooting).
In CSP missions, nodes on the network typically feature a command-line-based interface for configuration and debugging purposes [23]. This may allow someone with access to it to interact with the current node and other nodes on the network.
@@
CCSDS Space Communication Protocols. The CCSDS Space Communication protocols are a vast set of standardized protocols used for different purposes in space communications. A major CCSDS protocol relevant for SmallSat TM and TC is called spacepacket. This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) to define how TCs and TM are encoded and transported. The PUS defines services (and thus sets of TC/TM) for functionality that satellite missions likely require, including large data transfer or event reporting [59]. Moreover, mission designers can tailor the PUS standard by selecting a subset of services and sub-services relevant to their mission needs. Furthermore, it is possible to define custom Services or implement numerous custom functionalities.
@@
Proprietary Protocols. In addition to these widespread ecosystems, vendor-specific sets of protocols may be usable on top of more standard protocols or transport higher-layer packets of known protocols. Some basic small satellites may also have entirely ad hoc protocols that are limited to use with a single mission or a set of satellites by a specific institution. These proprietary protocols are not ideal candidates for constructing honeypots as an ecosystem of one mission that is unique to a small set of missions will look out of place for other missions, preventing generalization.
@@
The choice of ecosystem also influences the choice of GSS as an MCS usually focuses on a single protocol stack. In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services. These protocols include Telnet, SSH, VNC, FTP, and HTTP. Telnet, SSH, and VNC are used to remotely connect to the MOC workstations. HTTP is used for web interfaces that operators use to visualize mission data, and FTP is used to download TM data from the mission data repository. If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
2.6 Space Systems’ Tactics, Techniques and Procedures (TTPs)
@@
Tactics, Techniques, and Procedures (TTPs) describe the behavior of a malicious actor in a structured scheme to understand how they might execute an attack [56, 47]. Several frameworks have been introduced to standardize space systems’ TTPs. There are two MITRE-style frameworks, the SPARTA matrix [68] and the SPACE-SHIELD matrix [4]. These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise.
3 Threat Model
@@
Following Fig. 1, we assume an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first. To gain initial access to the ground segment, an adversary needs to connect via one of exposed network protocols depicted in Sec. 2.5, namely, VNC, Telnet, TCP/IP, etc., which correspond with the operational protocols used in real satellite missions.
From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission. Finally,
in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work.
Figure 2: HoneySat framework high-level architecture design.
4 HoneySat Framework Design
@@
In this section, we explain the objectives we aim to achieve (Sec. 4.1), the design principles we follow to meet such objectives (Sec. 4.2), and the overall design of our Space  (Sec. 4.4), and Ground (Sec. 4.3) segments.
4.1 HoneySat’s Design Objectives
@@ 1. DO-1
    Capability to Capture Rich Interaction Data. As explained in Sec. 1, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs. As such, our first objective for HoneySat is to have capability to capture rich interaction data.
@@ 2. DO-2
    Provide Deception. As we discussed in Sec. 1, honeypots’ nature must remain covert to entice adversaries into interacting with it. As such, our second objective is for HoneySat’s nature to remain hidden from adversaries.
@@ 3. DO-3
    Provide Extensibility and Customizability. A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs, in HoneySat’s case, we must be able to support multiple SmallSats. For example, a particular SmallSat may use CSP or CCSDS ecosystems. Additionally, each SmallSat mission has a particular orbit that must be customizable. As such, our third objective is for HoneySat to provide extensibility and customizability.
4.2 HoneySat’s Design Principles
@@ 1. DP-1
    High-Interaction Simulation. As we described in Sec. 2.1, high-interaction honeypots give adversaries the same or almost identical interaction opportunities as a real satellite. Thus, they can log many of the TTPs discussed in 2.6. For these reasons, we selected high-interaction simulation as our first design principle. This design principle is based on design objective DO-1.
@@ 2. DP-2
    Realistic Simulation. As we discussed in Sec. 2.2, satellites keep track of their orbit location, the position of the sun, among others. These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system. This design principle is based on design objective DO-2.
@@ 3. DP-3
    Modularity. As we explained in Sec. 2, satellites are complex and diverse systems. There is no one-size-fits-all solution. To tackle this problem, we designed HoneySat to be modular. Modularity allows us to insert, remove, or change any part of our honeypot framework without implementing a new honeypot from scratch. This design principle is based on our objective DO-3.
We now describe HoneySat’s architecture design and how this architecture integrates the design principles described above. At the highest level, our framework’s architecture, depicted in Fig. 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations. Table 2 illustrates how HoneySat’s simulation components match the components of a real satellite mission.
@@
Table 2: Correspondence between real satellite mission components and HoneySat’s design components.
4.3 Ground Segment Design
@@
Following Sec. 2.3, the purpose of the HoneySat’s ground segment is to simulate the ground segment assets, e.g., the ground segment software. To accomplish this, our design includes the following components: 1) the Exposed Network Protocols, 2) the Ground Segment Software, 3) the Radio Simulator, 4) the Ground Personality, and 5) the Logging Repository.
@@
1) Exposed Network Protocols. To provide adversaries with feasible access to our honeypot, HoneySat exposes multiple means of interaction over a network such as the Internet. We designed HoneySat to support four different interaction methods using different protocols, namely VNC, Telnet, Web, and Access to the ground station via raw packet transmitting capability (following design principle DP-3).
@@
2) Ground Software. As discussed in Sec. 2.3, the ground software includes mission control software (MCS) and ground station control software (GSCS). We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation (DP-1).
@@
3) Radio Simulator. As discussed in Sec. 2.3, communication between the satellite and the ground station is possible only during a pass. The purpose of the Radio Simulator is to control communication between the ground and space segment simulations to simulate real orbital passes. The Radio Simulator design follows design principles DP-2 and DP-3.
@@
4) Ground Configuration. The ground configuration is a series of settings for the various ground segment-specific configurations, e.g., the satellite mission logo in the Web Interface.
@@
5) Logging Repository. This repository logs data such as the TM/TC traffic to and from the ground station software and the logging attempts received in the web interface. We designed the ground segment simulation logs to be categorized and timestamped. The logging design follows design principle DP-3.
4.4 Space Segment Design
@@
The purpose of the HoneySat’s space segment is to mimic the spacecraft, as discussed in Sec. 2.3. To accomplish this, our design includes the following components: 1) the satellite’s Flight Software Runtime, 2) the Satellite Simulator, 3) the Flight Software Services, 4) the Satellite Personality, and 5) the Logging Repository.
@@
1) Flight Software Runtime. As discussed in Sec. 2.4, the satellite flight software manages all critical functions required for the mission operation, such as interacting with hardware peripherals, processing TCs, and sending TM. For this to work we need an environment where services that handle TC intended to run on flight software can be run. We reuse existing compatibility or testing wrappers to run the relevant parts of flight software for HoneySat to provide nearly identical interactions to an adversary. This produces a high-interaction honeypot simulation that follows design principle DP-1.
@@
2) Satellite Simulator. One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know. The Satellite Simulator is designed to solve this problem. It simulates all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function. The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripheral’s data. For example, suppose a particular flight software was implemented to utilize a specific lithium-ion battery pack. In that case, the Satellite Simulator’s power system simulation simulation can simulate the basic behavior of the battery. We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled, following our DP-3.
@@
3) Flight Software Services. The flight software services that process TC and provide TM act as the bridge between the Satellite Simulator and the rest of the Honeypot. In case a service requires some data from a subsystem of the satellite or a command is sent to a subsystem, it is passed to the satellite simulator instead. The satellite state module routes the service’s messages to and from the Satellite Simulator. It uses a list of message IDs that can be customized based on the protocol ecosystem and software architecture.
@@
4) Satellite Personality. The satellite personality module is designed to provide a central configuration location for the space segment. It includes FS and Satellite Simulator configuration values such as the satellite’s battery capacity. The satellite personality allows our framework to be easily customized.
@@ 4.5 Theory of Operation
An adversary gets initial access via one of the Exposed Network Protocols ①. On interaction with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②. The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker. The attacker can then interact with the Ground Software, while their actions are reported to the Log Database. They might want to try to gain more privileges on the ground segment or try to send TC to the space segment. Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator ③ to handle all simulated radio frequency communications.
When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location. If so, the Radio Simulator forwards the TC to the Flight Software Runtime ④. When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤. In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator for this ⑥. The Satellite Personality contains the configuration used by the Satellite Simulator. Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Log Database.
5 HoneySat’s Implementation
Having laid out HoneySat’s design we now explain how we implemented the Satellite Simulator (Sec. 5.1), the generic CSP mission honeypot (Sec. 5.2), the ground segment (Sec. 5.2.1), and the space segment (Sec. 5.2.2).
5.1 Satellite Simulator Implementation

### v2

@@ II-E Small Satellite Protocol Ecosystems
SmallSat missions can often be categorized by the adoption of protocols and their corresponding philosophies. Currently, there are two major protocol ecosystems which SmallSat missions can adopt: CSP and CCSDS.
@@
Cubesat Space Protocol. The Cubesat Space Protocol (CSP) family of protocols is a one-stop solution for SmallSat missions [55]. CSP is implemented as an open-source C library called libCSP [55] and follows the TCP/IP model, including transport and routing protocols and multiple layer 2 interfaces such as I2C (Inter-Integrated Circuit), and ZeroMQ (ZMQ) for transmission on TCP/IP networks [56].
In the context of a satellite mission, CSP connects the ground segment and satellite subsystems as part of a CSP network where each subsystem is identified as a node. Sending a TC is as simple as sending a CSP packet with the address of the corresponding subsystem node.
@@
In case the satellite is not currently performing a pass, the packet is dropped. There are also default services running on standardized ports that are enabled by default when building libCSP [55], for example for network diagnostic (ping) or basic operations (rebooting).
The straightforwardness of CSP makes it particularly relevant for the honeypot use case as the internal structures of any CSP-based mission will look similar, so distinguishing two missions may be difficult.
@@
CCSDS Space Communication Protocols. The Consultative Committee for Space Data Systems (CCSDS) Space Communication protocols are a set of standardized protocols used for different purposes in space communications [57]. A relevant CCSDS protocol for SmallSat TM and TC is the Space Packet protocol. This protocol is used in combination with the ECSS Packet Utilization Standard (PUS) [58] to define how TCs and TM are encoded and transported. PUS defines services (and thus sets of TC/TM) for functionality that satellite missions require, including large data transfer or event reporting [57].
@@
Proprietary Protocols. In addition to CSP and CCSDS, vendor-specific protocols may be used on top of more standard protocols. Some basic small satellites may have entirely ad-hoc protocols used on a single mission, for example the ESTCube-1 uses a custom protocol [59]. These proprietary protocols are not ideal candidates for constructing honeypots as the ecosystem of one mission will look out of place for other missions, preventing generalization.
@@
The choice of ecosystem also influences the choice of GSS as an MCS usually supports a single protocol stack [45]. In addition to the space protocol ecosystems above, satellite operators use well-known network protocols to connect with the mission’s ground segment services. These protocols include Telnet, SSH, VNC, FTP, and HTTP. Telnet, SSH, and VNC are used to remotely connect to the MOC workstations [11]. HTTP is used for web interfaces that operators use to visualize mission data. FTP is used to download TM data from the mission data repository. If an adversary were to compromise any of these services, they could use it as a stepping stone to target the satellite itself.
II-F Space Systems’ Tactics, Techniques and Procedures (TTPs)
@@
Tactics, Techniques, and Procedures (TTPs) describe a malicious actor’s behavior in a structured scheme to understand how they might execute an attack [60, 61]. Several frameworks have been introduced to standardize space systems’ TTPs. There are two MITRE-style frameworks, the SPARTA matrix [62] and the SPACE-SHIELD matrix [63]. These matrices list and describe space security-specific tactics and techniques such as initial access and ground segment compromise. For the remainder of this paper we use the SPACE-SHIELD matrix to analyze HoneySat.
III Threat Model
@@
Following Fig. 2, we assume that an adversary willing to compromise a satellite can only interact with the space segment simulation by sending TCs from the ground segment first. To gain initial access to the ground segment, an adversary needs to connect via one of the exposed network protocols discussed in Sec. II-E, namely, VNC, Telnet, TCP/IP, etc., which correspond to the operational protocols used in real satellite missions. From there, an adversary may try to launch different commands to take full control and/or compromise the services offered by the satellite’s mission as depicted in Fig. 3. Finally, in this paper, the modeling of physical radio communication between the space and ground segments, which is commonly used in practice, is considered out of scope and left for future work. The threat model is further referenced in Sec. IV-E.
IV HoneySat Framework Design
@@
In this section, we explain the objectives we aim to achieve (Sec. IV-A), the design principles we follow to meet such objectives (Sec. IV-B), and the overall design of our Space  (Sec. IV-D), and Ground (Sec. IV-C) segments.
IV-A HoneySat’s Design Objectives
@@ 1. DO-1
    Capability to Capture Rich Interaction Data. As explained in Sec. I, the number one objective of any honeypot is to capture interaction data from which we can derive knowledge on adversaries’ TTPs. As such, our first objective for HoneySat is to have the capability to capture rich interaction data.
@@ 2. DO-2
    Provide Deception. As we discussed in Sec. I, honeypots’ nature must remain covert to entice adversaries to interact with it. As such, our second objective is for HoneySat’s nature to remain hidden from adversaries.
@@ 3. DO-3
    Provide Extensibility and Customizability. A framework’s main purpose is to provide generic functionality that can be customized to meet the user’s needs. In HoneySat’s case, we must be able to support multiple SmallSats. For example, a particular SmallSat may use CSP or CCSDS ecosystems. Additionally, each SmallSat mission has a particular orbit that must be customizable. As such, our third objective is for HoneySat to provide extensibility and customizability.
Figure 2: HoneySat framework high-level architecture design.
IV-B HoneySat’s Design Principles
@@ 1. DP-1
    High-Interaction Simulation. As we described in Sec. II-A, high-interaction honeypots give adversaries the same or almost identical interaction opportunities as a real satellite. For these reasons, we selected high-interaction simulation as our first design principle. This design principle is based on design objective DO-1.
@@ 2. DP-2
    Realistic Simulation. As we discussed in Sec. II-B, satellites keep track of their orbit location, the position of the sun, among other details. These details must be simulated by our honeypot; otherwise, they may alert adversaries that they are interacting with a fake system. For this reason, we selected realistic simulation as our second design principle. This design principle is based on design objective DO-2.
@@ 3. DP-3
    Modularity. As we explained in Sec. II, satellites are complex and diverse systems. There is no one-size-fits-all solution. To tackle this problem, we designed HoneySat to be modular. Modularity allows us to insert, remove, or change any part of our honeypot framework without implementing a new honeypot from scratch. This design principle is based on our objective DO-3.
We now describe HoneySat’s architecture design and how this architecture integrates the design principles described above. At the highest level, our framework’s architecture, depicted in Fig. 2, consists of two sets of simulations, the space segment simulations and the ground segment simulations. Table II illustrates how HoneySat’s simulation components match the components of a real satellite mission.
TABLE II: Correspondence between real satellite mission components and HoneySat’s design components.
@@
IV-C Ground Segment Design
@@
Following Sec. II-C, the purpose of the HoneySat’s ground segment (depicted in Fig. IV) is to simulate the ground segment assets, e.g., the ground segment software. To accomplish this, our design includes the following components: 1) the Exposed Network Protocols, 2) the Ground Segment Software, 3) the Radio Simulator, 4) the Ground Personality, and 5) the Logging Repository.
@@
1) Exposed Network Protocols. To provide adversaries with feasible access to our honeypot, HoneySat exposes multiple means of interaction over a network such as the Internet. We designed HoneySat to support four different interaction methods using different protocols, namely VNC, Telnet, Web, and Access to the ground station via raw packet transmitting capability; following the design principle DP-3. We selected these protocols based on data obtained from our satellite operator survey, discussed in Sec. VI-C, which revealed that satellite operators do use remote access tools, such as web interfaces and screen sharing, to operate satellites.
@@
2) Ground Software. As discussed in Sec. II-C, the ground software includes mission control software (MCS) and ground station control software (GSCS). We leverage existing ground software, allowing HoneySat to provide a high-interaction simulation; thereby following design principle DP-1.
@@
3) Radio Simulator. As discussed in Sec. II-C, communication between the satellite and the ground station is possible only during a pass. The purpose of the Radio Simulator is to mimic real orbital passes by enabling and disabling communication between HoneySat’s ground and space segment simulations at the appropriate times. The Radio Simulator design follows the design principles DP-2 and DP-3.
@@
4) Ground Configuration. The ground configuration is a series of settings for the ground segment-specific configurations, e.g., the satellite mission logo in the Web Interface.
@@
5) Logging Repository. This repository records data such as the TM/TC traffic to and from the ground station software and the logging attempts received in the web interface. We designed the ground segment simulation logs to be categorized and timestamped. The logging repository design follows design principle DP-3.
IV-D Space Segment Design
@@
The purpose of the HoneySat’s space segment is to mimic the spacecraft, as discussed in Sec. II-C. To accomplish this, our design includes the following components: 1) the satellite’s Flight Software Runtime, 2) the Satellite Simulator, 3) the Flight Software Services, 4) the Satellite Personality, and 5) the Logging Repository.
@@
1) Flight Software Runtime. As discussed in Sec. II-D, the satellite flight software manages all critical functions required for the mission operation, such as interacting with hardware peripherals, processing TCs, and sending TM. For this to work we need an environment where services that handle TC intended to run on flight software can be run. We reuse existing compatibility or testing wrappers to run the relevant parts of flight software for HoneySat to provide nearly identical interactions to an adversary. This produces a high-interaction honeypot simulation that follows design principle DP-1.
@@
2) Satellite Simulator. One of the biggest challenges when designing HoneySat was simulating all the physical processes, e.g., attitude, that satellites need to know. The Satellite Simulator is designed to solve this problem by simulating all the necessary satellite subsystems, e.g., the Electrical Power System, and the sensors, e.g., a GPS receiver, that a satellite requires to function. The purpose of the Satellite Simulator is to abstract away the various hardware peripherals that communicate with the flight software by simulating these peripherals’ data. For example, suppose that a particular flight software was implemented to utilize a specific lithium-ion battery pack. In that case, the Satellite Simulator’s power system simulation can simulate the basic behavior of the battery. We designed the Satellite Simulator to be modular so that new subsystems and sensors can be easily added, modified, or disabled according to our design principle DP-3.
@@
3) Flight Software Services. The flight software services that process TC and provide TM act as the bridge between the Satellite Simulator and the rest of the honeypot. In case a service requires some data from a subsystem of the satellite or a command is sent to a subsystem, it is passed to the satellite simulator instead. The satellite state module routes the service’s messages to and from the Satellite Simulator. It uses a list of message IDs that can be customized based on the protocol ecosystem and software architecture complying with our design principles DP-1 and DP-3.
@@
4) Satellite Personality. The satellite personality module is designed to provide a central configuration location for the space segment. It includes FS and Satellite Simulator configuration values such as the satellite’s battery capacity. The satellite personality makes our framework easily customizable; thus following our design principle DP-3.
@@ IV-E Theory of Operation
An adversary gets initial access via one of the Exposed Network Protocols ①. On interacting with the Simulated Ground Segment, an attacker is presented with access to the Ground Software ②. The configuration of this ground software is defined by the Ground Configuration, which allows the ground software to look like one of many different missions from its protocol ecosystem to the attacker. The attacker can then interact with the Ground Software, while their actions are reported to the Logging Repository. They might want to try to gain more privileges on the ground segment or try to send TC to the space segment. Instead of deploying a ground station with an RF transmitter and associated hardware, we employ the Radio Simulator ③ to handle all simulated radio frequency communications.
When an attacker uses the ground segment to send a valid TC or raw packets, the Radio Simulator checks whether the satellite configured in the Satellite Personality is currently passing over the designated ground station location. If so, the Radio Simulator forwards the TC to the Flight Software Runtime ④. When the Flight Software Runtime receives a command, it will run the corresponding Flight Software Service to compute a response ⑤. In case it requires either changes to an on-board system’s state or a sensor value to execute the TC, it will invoke the Satellite Simulator ⑥. The Satellite Personality contains the configuration used by the Satellite Simulator. Once a TC is processed, the resulting TM is routed back through the Radio Simulator to either the attacker or the ground software used by the attacker ⑦-⑫, and is also recorded in our Logging Repository.
V HoneySat’s Implementation
Having laid out HoneySat’s design we now explain how we implemented the Satellite Simulator (Sec. V-A), the generic CSP mission honeypot (Sec. V-B), the ground segment (Sec. V-B1), and the space segment (Sec. V-B2).
V-A Satellite Simulator Implementation